中級技術英語 1 単位 (選択)

Technical English  カーペンターウォルター · 講師/ 機械工学科生産システム講座

【授業目的】 The primary aim of the course is to enhance reading and writing skills

in technical–especially scientific and engineering–English. 【授業概要】The primary focus will be the development of three English skills:

listening comprehension, reading and writing. The key objective will be to increase and broaden knowledge of technical and scientific vocabulary. As a result, students should be better able to understand and produce English technical writing, especially in the field of Engineering. Therefore, the emphasis of the course content is intended to strengthen and expand both comprehension and expressive abilities. It is anticipated that this will include improving the basic aspects of clear writing: important aspects of grammar, vocabulary, and recurring patterns of typical sentences commonly found in technical English.

【キーワード】 technical English， scientific technical articles， technical term 【先行科目】『初級技術英語』 (1.0) 【関連科目】『上級技術英語』 (1.0) 【履修要件】 You should have gotten the credit of ”Basic Technical English.” 【履修上の注意】日亜 STC 学生は必ず受講すること．日亜 STC 学生でない学生が

この科目を受講する場合は自分の学科のコーディネータに連絡し受講許可を受けること．

【到達目標】

1. To master vocabulary and sentence formation commonly encountered in scientific technical articles, journals and books.

2. To develop three English skills: listening comprehension, reading and writing. 【授業計画】

1. Conventions of English 1

2. Conventions of English 2

3. Vocabulary: correct spelling

4. Vocabulary: how to communicate to an English-speaking person the spelling of a word or phrase, homonyms,synonyms and commonly misspelled or misunderstood English words

5. Noun and verb agreement 1

6. Noun and verb agreement 2

7. Reading skills 1

8. Reading skills 2 9. Differences between academic and professional writing 1

10. Differences between academic and professional writing 2

11. Important points to avoid in technical writing 1

12. Important points to avoid in technical writing 2

13. How to describe commonly used adjectives, phrases and words that describe location, phrases and words that describe the frequency of occurrence or the unique qualities of a thing, a process, or research.

14. How to understand commonly used adjectives, phrases and words that describe location, phrases and words that describe the frequency of occurrence or the unique qualities of a thing, a process, or research.

15. When and how to compose formal communication: letters of acceptance, letters of inquiry, project proposal letters.

16. Electronic forms of communication (E-mail), using the internet to conduct research, how to cite internet sources of information, what to avoid when using the internet for research purposes.

【成績評価基準】 The overall grade will be based on the mid-term and final examinations

and homework assignments. 【学習教育目標との関連】 [JABEE 関連] 【教科書】 Presenting Science: A Technical English Course for Engineering Majors.

(Second Edition). 2008. Macmillan Language House. 【参考書】 [参考資料] 【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216144

【対象学生】Nichia STC students and a limitted number of students who are

permitted to get this subject. 【連絡先】

⇒ 西口 (創成学習開発センター 3 階, 088-656-7304, nishiguchi@cicee.tokushi

ma-u.ac.jp) ���� (オフィスアワー: 水曜日 12時-1時) ⇒ 橋爪 (E 棟 3 階南 D-2, 088-656-7473, tume@ee.tokushima-u.ac.jp) ����

Technical English 1 unit (selection)

Walter Carpenter · Associate Professor / Production Systems Engineering, Department of Mechanical Engineering

Target⟩ The primary aim of the course is to enhance reading and writing skills in

technical–especially scientific and engineering–English. Outline⟩ The primary focus will be the development of three English skills:

listening comprehension, reading and writing. The key objective will be to increase and broaden knowledge of technical and scientific vocabulary. As a result, students should be better able to understand and produce English technical writing, especially in the field of Engineering. Therefore, the emphasis of the course content is intended to strengthen and expand both comprehension and expressive abilities. It is anticipated that this will include improving the basic aspects of clear writing: important aspects of grammar, vocabulary, and recurring patterns of typical sentences commonly found in technical English.

�������⟩ technical English, scientific technical articles, technical term Fundamental Lecture⟩ “Basic Technical English”(1.0) Relational Lecture⟩ “Advanced Technical English”(1.0) Requirement⟩ You should have gotten the credit of ”Basic Technical English.” Notice⟩ This subject is a sompulsory one for Nichia STC students. If you are not

Nichia STC students and want to attend this class, please contact a coordinator in your department for getting permission of your attendance.

Goal⟩

1. To master vocabulary and sentence formation commonly encountered in scientific technical articles, journals and books.

2. To develop three English skills: listening comprehension, reading and writing.

Schedule⟩

1. Conventions of English 1

2. Conventions of English 2

3. Vocabulary: correct spelling

4. Vocabulary: how to communicate to an English-speaking person the spelling of a word or phrase, homonyms,synonyms and commonly misspelled or misunderstood English words

5. Noun and verb agreement 1

6. Noun and verb agreement 2

7. Reading skills 1

8. Reading skills 2 9. Differences between academic and professional writing 1

10. Differences between academic and professional writing 2

11. Important points to avoid in technical writing 1

12. Important points to avoid in technical writing 2

13. How to describe commonly used adjectives, phrases and words that describe location, phrases and words that describe the frequency of occurrence or the unique qualities of a thing, a process, or research.

14. How to understand commonly used adjectives, phrases and words that describe location, phrases and words that describe the frequency of occurrence or the unique qualities of a thing, a process, or research.

15. When and how to compose formal communication: letters of acceptance, letters of inquiry, project proposal letters.

16. Electronic forms of communication (E-mail), using the internet to conduct research, how to cite internet sources of information, what to avoid when using the internet for research purposes.

Evaluation Criteria⟩ The overall grade will be based on the mid-term and final

examinations and homework assignments. Relation to Goal⟩ [JABEE 関連] Textbook⟩ Presenting Science: A Technical English Course for Engineering

Majors. (Second Edition). 2008. Macmillan Language House. Reference⟩ [参考資料] Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216144

Student⟩ Nichia STC students and a limitted number of students who are permitted

to get this subject.

Contact⟩

⇒ Nishiguchi (Innovation Center 3rd floor, +81-88-656-7304, nishiguchi@cice

e.tokushima-u.ac.jp) ���� (Office Hour: Wed 12pm-1pm) ⇒ Hashizume (E 棟 3 階南 D-2, +81-88-656-7473, tume@ee.tokushima-u.ac.j

p) ����